OpenStudy (anonymous):

I understand dot products, but this question is confusing me: Use the component form of dot product of two arbitrary vectors u and w to prove the trigonometric identity cos (θ- β) = cos θ cos β + sin θ sin β

OpenStudy (anonymous):

the component for of a vector A is written \[A = A _{x}i + A _{y}j\] A dot B = \[A _{x}B _{x} + A _{y}B _{y}\] Can you do the rest?

OpenStudy (anonymous):

I don't understand what A and what B are supposed to be in his particular problem

OpenStudy (anonymous):

is this the one you wanted me to look at?

OpenStudy (anonymous):

Yes it is

OpenStudy (anonymous):

@amistre64

OpenStudy (amistre64):

found it :)

OpenStudy (anonymous):

Awesome, let me know if you have any ideas

OpenStudy (amistre64):

A and B were generalities what is the cos of an angle in dot product format?

OpenStudy (anonymous):

The cos of the angle is usually used just to get one of the vectors in line with the other

OpenStudy (amistre64):

\[cos(a)=\frac{\vec u\cdot \vec v}{|\vec u|~|\vec v|}\]

OpenStudy (amistre64):

now, it looks like cos(a)cos(b) + sin(a)sin(b) is a dot product of: (cos(a),sin(a)) and (cos(b),sin(b))

OpenStudy (amistre64):

cos(a-b) has a trig identity: it is equal to cos(a)cos(b) + sin(a)sin(b)

OpenStudy (amistre64):

you are trying to prove this with a dotproduct method

OpenStudy (amistre64):